# Disentangling quantum autoencoder

### v1

@@ Abstract
Entangled quantum states are highly sensitive to noise, which makes it difficult to transfer them over noisy quantum channels or to store them in quantum memory. Here, we propose the disentangling quantum autoencoder (DQAE) to encode entangled states into single-qubit product states. The DQAE provides an exponential improvement in the number of copies needed to transport entangled states across qubit-loss or leakage channels compared to unencoded states. The DQAE can be trained in an unsupervised manner from entangled quantum data. We demonstrate variational quantum algorithms for general states, and a Metropolis algorithm for stabilizer states. For particular classes of states, the number of training data needed to generalize is surprisingly low: For stabilizer states, DQAE generalizes by learning from a number of training data that scales linearly with the number of qubits, while only $1$ training data is sufficient for states evolved with the transverse-field Ising Hamiltonian. Our work provides practical applications for enhancing near-term quantum computers.
@@ I Introduction
Entanglement is a key feature of quantum mechanics that describes the non-local quantum correlations, and is a necessary ingredient for achieving quantum advantages in quantum communication, quantum sensing and quantum computing [1]. However, at the same time, entanglement makes quantum states fragile, as the non-local correlations are highly susceptible to noise that affects quantum states in experiment. In particular, even noise that acts only on a local subsystem of the entangled state can already collapse the global wavefunction into a classical mixture of states without useful quantum correlations.
@@
To make states robust against noise, various methods have been proposed, such as quantum error correction where the logical information is encoded redundantly into many physical qubits [2]. However, quantum error correction is highly resource-intensive, and not possible in noisy intermediate-scale quantum computers [3]. Furthermore, error correction may be infeasible in scenarios where physical constraints exist, such as a limited number of qubits or long-distance communication between parties.
@@
Alternatively, can reducing the entanglement of states enhance their robustness to noise?? In classical settings, reducing non-local correlations to simplify data has proven to be a fruitful approach. For example, in classical neural networks, variational autoencoders have been used to disentangle highly correlated information into factorized probability distributions [4, 5]. This has allowed to find simpler representations of complicated data that can enhance machine learning.
@@
In quantum information, quantum autoencoders have been proposed to compress high-dimensional quantum states into a lower-dimensional subspace [6, 7, 8]. They can finding a more compact representation of quantum states [9, 10, 11, 12] to enhance information processing. Notably, quantum autoencoders can be used to denoise quantum states [13, 14, 15], and discover new quantum error correction codes [16]. They have been applied to enhance quantum communication in photonic systems [17]. Quantum autoencoders have been fruitful in experiment, with several experimental demonstrations [18, 15, 17].
@@
Quantum autoencoders have been implemented using variational quantum algorithms [19], in which the unitary is represented by a parameterized quantum circuit that is trained on a few example states. To train the circuit parameters, the encoding unitary is learned [20, 21]. The training can be challenging, requiring one to avoid exponential small gradients in barren plateaus [22]. Further, to successfully converge to a good solution, one often requires Overparameterization [23, 24, 25, 26, 27, 28]. Additionally, one would like to generalize, i.e. that from a few training states drawn from some distribution, the trained circuit is able to perform well on arbitrary states drawn from the distribution [29, 30, 31, 32, 33, 28].
@@
However, the literature on quantum autoencoders mainly focuses on reducing the number of qubits, i.e. finding a lower-dimensional representation of quantum states. Compressing information to less qubits implies that quantum information is encoded into non-local correlations, strongly increasing the amount of entanglement. Thus, the current types of quantum autoencoder produces highly entangled states, which are sensitive to noise, and have a complicated representation.
@@
Here, we propose the disentangling quantum autoencoder (DQAE). It transforms entangled $N$-qubit quantum states into a tensor products of $N$ single-qubit quantum states, removing entanglement and factorizing the quantum correlations. We apply the DQAE for the task of transporting entangled state through a qubit-loss channel. In comparison to unencoded transport, the DQAE has an exponential advantage in the number of copies needed to transfer at least one error-free state. The DQAE can be trained in an unsupervised manner using entangled quantum data via variational quantum algorithms. We show that the gradient can be computed efficiently from simple Pauli measurements or Bell measurements. We show that for structured circuit types, barren plateaus can be absent, allowing for efficient training. We characterize the generalization capability, i.e. whether training on a few states allows the DQAE to disentangle unseen test states, using the data quantum Fisher information metric (DQFIM). For Ising-type evolution, we astonishingly require only $L=1$ training states and $M\propto N$ circuit parameters to generalize. For stabilizer states, we propose a Metropolis algorithm to train a Clifford ansatz that generalizes using only $L\propto N/4$ training states. With our results, quantum information processing with near-term quantum computers can be greatly enhanced.
@@ II Model
The DQAE $U$, as shown in Fig. 1, disentangles a set $W$ of $N$-qubit quantum state $|\psi\rangle\in W$ into a tensor product of $N$ single-qubit states
@@
where $|\phi_{j}\rangle$ are single-qubit states that depend on $|\psi\rangle$. As $U$ is unitary, one can reconstruct the entangled original state via the inverse $U^{\dagger}$
@@
From simple counting arguments, one can see that a single DQAE $U$ cannot compress the set of all possible $N$-qubit states $W=\mathcal{S}(2^{N})$ into single-qubit states at the same time. In particular, to represent any $N$-qubit state $|\psi\rangle$ as single-qubit tensor products, one requires $O(2^{N})$ single-qubit states [34, 35, 36]. However, efficient compression is possible when we consider the DQAE for compressing a restricted set $W$ of states. We note that while we only consider compression to $N$ qubits, one can also reduce the number of single-qubit product states to $K\leq N$. Here, one can further enhance the capacity of the DQAE by including classical information into the $N-K$ qubits that were reduced as shown in Appendix A.
@@
Figure 1: Disentangling quantum autoencoder (DQAE) disentangles $N$-qubit state $|\psi\rangle$ into a tensor product of $N$ single-qubit states $|\phi_{j}\rangle$ via unitary $U(\boldsymbol{\theta})$, and can recover the original state again via the inverse $U^{\dagger}(\boldsymbol{\theta})$. The DQAE can be trained in an unsupervised manner from entangled data. For transferring quantum states via a qubit-loss channel, DQAE gives an exponential improvement in number of copies needed.
@@
Figure 2: Disentangling quantum autoencoder (DQAE) to transfer quantum states across qubit loss channels with exponential advantage. Entangled input $N$-qubit state $|\psi\rangle$ is transformed by $U(\boldsymbol{\theta})$ into $\bigotimes_{j=1}^{N}|\phi_{j}\rangle=U|\psi\rangle$,a tensor product of $N$ single-qubit states $|\phi_{j}\rangle$. The transformed state is then sent through a qubit loss channel $\mathcal{E}$, and stored in a quantum register by the receiver. We repeat this process $R$ copies of $|\psi\rangle$, and finally store $NR$ qubits in the quantum register. Some qubits are lost due to the channel $\mathcal{E}$ (red crosses), which we detected using the leakage detection protocol of Fig. 3. At least one of each single-qubit state survives with high probability when $R\propto\log(N)$. In this case, we collect $k=1,\dots,N$ single-qubit states $|\phi_{j}\rangle$ unaffected by loss (green circle). With entangler $U^{\dagger}$ we then reconstruct the original state $|\psi\rangle=U^{\dagger}(\boldsymbol{\theta})\bigotimes_{j=1}^{N}|\phi_{j}\rangle$.
@@ III DQAE for qubit-loss channels
[…]
$\mathcal{E}_{k}(\rho)=(1-q)\rho+q\text{tr}_{\bar{k}}(\rho)\otimes|\bot\rangle\langle\bot|\,,$ (3)
where with probability $q$ is lost. Here, $\text{tr}_{\bar{k}}(\rho)$ traces out the $k$th qubit. We assume loss occurs on all qubits equally with the total channel
@@
For a fully entangled state, loss on any qubit collapses the state into a mixture, and it cannot be used for further processing. As simple example, let us consider the $N$-qubit GHZ state
@@
Loss of any qubit collapses the state into a $N-1$ qubit mixed state
@@
the channel only traces out the $k$th qubit state $|\phi_{k}\rangle$, while the other $N-1$ qubits remain in a pure state. Thus, product states are more robust, which we will showcase in the following.
@@ III.1 Unencoded transfer
First, let us consider a direct unencoded transfer of the $R$ copies. We apply channel $\mathcal{E}(|\psi\rangle)$ on the pure $N$-qubit entangled state $|\psi\rangle$ and with a chance of $(1-q)^{N}$ no loss occurs on any qubit. The chance of successfully transferring the state has an exponentially small probability. Now we repeat the protocol with $R$ copies of $|\psi\rangle$. The probability of at least one copy successfully being transferred has a failure probability of
@@
which implies that we require exponentially large $R$ with $N$ for fixed error probability $Q_{\text{e}}$.
@@ III.2 DQAE
Now, we compress the incoming $N$-qubit state $|\psi\rangle$ into a single-qubit product state using the DQAE. Then, we have the $N$ qubit product state $U|\psi\rangle=\otimes_{j=1}^{N}|\phi_{j}\rangle$, where $|\phi_{j}\rangle\in\mathcal{H}^{2}$ is a single qubit state. Product states are robust against loss, as in the event of loss the remaining qubits remain pure. Our protocol for product states is as follows: We have $R$ copies of $U|\psi\rangle$. After applying $\mathcal{E}(U|\psi\rangle)$, some qubits are lost, however we assume that of the single qubit states $|\phi_{j}\rangle$, at least one of each has survived. One can check whether a qubit has survived with the leakage detection protocol [37] as shown in Fig.3, which also has been realized in experiment [38].
@@
[…]
We now compute the failure probability $Q_{\text{p}}$ of the DQAE. For $N=1$, the probability that at least one qubit survives the channel is $1-q^{R}$. For $N$ qubits, the product protocol has a failure probability
@@ III.3 Fidelity
We now show how to lower bound the fidelity of the protocol of Fig. 2. We have entangled state $|\psi\rangle$ and then run our protocol of Fig. 2. Let us assume that during the protocol we send $N$ copies of disentangled $U|\psi\rangle$ through the loss channel. We assume that for the $k$th copy, we loose all qubits except the $k$th qubit, remaining $\rho_{k}=\text{tr}_{\bar{k}}(|\psi\rangle)$. We can detect which qubits survived using the leakage detection of Fig. 3, and use the reconstruction step to assemble the $N$ remaining qubits $\rho_{k}$ into a single state $\rho=\bigotimes_{k}\rho_{k}$. If the disentangler is optimal, we can perfectly reconstruct the original stat via $|\psi\rangle=U^{\dagger}\rho U$, as each $\rho_{k}$ is pure.
However, what if the disentangler is not working perfectly? An optimal disentangler has $E=\frac{1}{N}\sum_{k=1}^{N}E_{k}(U|\psi\rangle)=1$, i.e. the disentangled state is a tensor product of single-qubit states.
If the disentangler does not optimally disentangle states, it creates an entangled state with $U|\psi\rangle$ with $E<1$. In this case, for the reconstruction step of Fig. 2, tracing out for each individual qubit $k$ results in a mixed state $\rho_{k}=\text{tr}_{\bar{k}}(U|\psi\rangle)$. The reconstructed state is then given by $\rho_{\text{f}}=U^{\dagger}\bigotimes_{k=1}^{N}\text{tr}_{\bar{k}}(U|\psi\rangle)U$. We can compute the fidelity between the original state $|\psi\rangle$ and the reconstructed state as
@@
which is $F=1$ only when DQAE encoded state $U|\psi\rangle$ is a product state, and is bounded by $2^{-N}\leq F\leq 1$.

### v2

# thanks: Corresponding author

@@ Abstract
Entangled quantum states are highly sensitive to noise, which makes it difficult to transfer them over noisy quantum channels or to store them in quantum memory. Here, we propose the disentangling quantum autoencoder (DQAE) to encode entangled states into single-qubit product states. The DQAE provides an exponential improvement in the number of copies needed to transport entangled states across qubit-loss or leakage channels compared to unencoded states. The DQAE can be trained in an unsupervised manner from entangled quantum data. For general states, we train via variational quantum algorithms based on gradient descent with purity-based cost functions, while stabilizer states can be trained via a Metropolis algorithm. For particular classes of states, the number of training data needed to generalize is surprisingly low: For stabilizer states, DQAE generalizes by learning from a number of training data that scales linearly with the number of qubits, while only $1$ training sample is sufficient for states evolved with the transverse-field Ising Hamiltonian. Our work provides practical applications for enhancing near-term quantum computers.
@@ I Introduction
Entanglement is a key feature of quantum mechanics that describes the non-local quantum correlations, and is a necessary ingredient for achieving quantum advantages in quantum communication, quantum sensing and quantum computing [1]. However, at the same time, entanglement makes quantum states fragile, as the non-local correlations are highly susceptible to noise that affects quantum states in experiments. In particular, even noise that acts only on a local subsystem of the entangled state can already collapse the global wavefunction into a classical mixture of states without useful quantum correlations.
@@
To make states robust against noise, various methods have been proposed, such as quantum error correction, where the logical information is encoded redundantly into many physical qubits [2]. However, quantum error correction is highly resource-intensive and not possible in noisy intermediate-scale quantum computers [3]. Furthermore, error correction may be infeasible in scenarios where physical constraints exist, such as a limited number of qubits or long-distance communication between parties.
@@
Alternatively, can reducing the entanglement of states enhance their robustness to noise? In classical settings, reducing non-local correlations to simplify data has proven to be a fruitful approach. For example, in classical neural networks, variational autoencoders have been used to disentangle highly correlated information into factorized probability distributions [4, 5]. This has allowed for finding simpler representations of complicated data that can enhance machine learning.
@@
In quantum information, quantum autoencoders have been proposed to compress high-dimensional quantum states into a lower-dimensional subspace [6, 7, 8]. They can find a more compact representation of quantum states [9, 10, 11, 12] to enhance information processing. Notably, quantum autoencoders can be used to denoise quantum states [13, 14, 15], and discover new quantum error correction codes [16]. They have been applied to enhance quantum communication in photonic systems [17]. Quantum autoencoders have been fruitful in experiments, with several experimental demonstrations [18, 15, 17].
@@
Quantum autoencoders have been implemented using variational quantum algorithms [19], in which the unitary is represented by a parameterized quantum circuit that is trained on a few example states. To train the circuit parameters, the encoding unitary is learned [20, 21]. The training can be challenging, requiring one to avoid exponentially small gradients in barren plateaus [22]. Further, to successfully converge to a good solution, one often requires overparameterization [23, 24, 25, 26, 27, 28]. Additionally, one would like to generalize, i.e. from a few training states drawn from some distribution, the trained circuit is able to perform well on arbitrary states drawn from the distribution [29, 30, 31, 32, 33, 28].
@@
However, the literature on quantum autoencoders mainly focuses on reducing the number of qubits, i.e. finding a lower-dimensional representation of quantum states. Compressing information to fewer qubits implies that quantum information is encoded into non-local correlations, strongly increasing the amount of entanglement. Thus, the current types of quantum autoencoders produce highly entangled states, which are sensitive to noise, and have a complicated representation.
@@
Here, we propose the disentangling quantum autoencoder (DQAE). It transforms entangled $N$-qubit quantum states into a tensor product of $N$ single-qubit quantum states, removing entanglement and factorizing the quantum correlations. We apply the DQAE for the task of transporting entangled state through a qubit-loss channel. In comparison to unencoded transport, the DQAE has an exponential advantage in the number of copies needed to transfer at least one error-free state. The DQAE can be trained in an unsupervised manner using entangled quantum data via variational quantum algorithms. The corresponding cost function is either based on local Pauli observables or the local purity of the state. We show that the gradient can be measured efficiently via Pauli or Bell measurements. We show that for structured circuit types, barren plateaus can be absent, allowing for efficient training. We characterize the generalization capability, i.e. whether training on a few states allows the DQAE to disentangle unseen test states, using the data quantum Fisher information metric (DQFIM). For Ising-type evolution, we astonishingly require only $L=1$ training states and $M\propto N$ circuit parameters to generalize. For stabilizer states, we propose a Metropolis algorithm to train a Clifford ansatz that generalizes using only $L\propto N/4$ training states. With our results, quantum information processing with near-term quantum computers can be greatly enhanced.
@@ II Model
The DQAE unitary $U$, as shown in Fig. 1, disentangles a set $W$ of $N$-qubit quantum states $|\psi\rangle\in W$ into a tensor product of $N$ single-qubit states
@@
where $|\phi_{j}\rangle$ are single-qubit states that depend on $|\psi\rangle$. Details of how the unitary $U$ is implemented and trained for such sets are provided in Sec IV. As $U$ is unitary, one can reconstruct the original entangled state via the inverse $U^{\dagger}$
@@
From simple counting arguments, one can see that a single DQAE $U$ cannot compress the set of all possible $N$-qubit states $W=\mathcal{S}(2^{N})$ into single-qubit states at the same time. In particular, to represent any $N$-qubit state $|\psi\rangle$ as single-qubit tensor products, one requires $O(2^{N})$ single-qubit states [34, 35, 36]. However, efficient compression is possible when we consider the DQAE for compressing a restricted set $W$ of states. We note that while we only consider compression to $N$ qubits, one can also reduce the number of single-qubit product states to $K\leq N$. Here, one can further enhance the capacity of the DQAE by including classical information into the NKN-K qubits that were reduced as shown in Appendix A.
@@
Figure 1: Disentangling quantum autoencoder (DQAE) disentangles NN-qubit state $|\psi\rangle$ into a tensor product of $N$ single-qubit states $|\phi_{j}\rangle$ via unitary $U(\boldsymbol{\theta})$, and can recover the original state again via the inverse $U^{\dagger}(\boldsymbol{\theta})$. The DQAE can be trained in an unsupervised manner from entangled data. For transferring a quantum state via a qubit-loss channel, DQAE gives an exponential reduction in the number of copies $R$ of the state.
@@
Figure 2: Illustration of the product protocol: Disentangling quantum autoencoder (DQAE) to transfer quantum states across qubit loss channels with exponential advantage. Entangled input NN-qubit state $|\psi\rangle$ is transformed by $U(\boldsymbol{\theta})$ into $\bigotimes_{j=1}^{N}|\phi_{j}\rangle=U|\psi\rangle$, a tensor product of $N$ single-qubit states $|\phi_{j}\rangle$. The transformed state is then sent through a qubit loss channel $\mathcal{E}$, and stored in a quantum register by the receiver. We repeat this process for $R$ copies of $|\psi\rangle$, and finally store NRNR qubits in the quantum register. Some qubits are lost due to the channel $\mathcal{E}$ (red crosses), which we detected using the leakage detection protocol of Fig. 3. At least one of each single-qubit state survives with high probability when $R\propto\log(N)$. In this case, we collect $k=1,\dots,N$ single-qubit states $|\phi_{j}\rangle$ unaffected by loss (green circle). With entangler $U^{\dagger}$ we then reconstruct the original state $|\psi\rangle=U^{\dagger}(\boldsymbol{\theta})\bigotimes_{j=1}^{N}|\phi_{j}\rangle$.
@@ III DQAE for qubit-loss channels
We now propose the DQAE algorithm, which is highly tolerant to qubit loss channels. First, let us describe the transport problem that we aim to solve: We assume that we have $R$ copies of an $N$-qubit state $|\psi\rangle$. We can send the $R$ copies of $|\psi\rangle$ one by one through a noisy quantum channel $\mathcal{E}(|\psi\rangle)$. Now, we want to achieve that at least one copy passes through the channel and arrives as a pure state. We assume that each qubit is subject to a qubit loss channel or leakage error [37], where the $k$th qubit is transferred to an inaccessible leakage state $|\bot\rangle$ orthogonal to the other qubit states.
$\mathcal{E}_{k}(\rho)=(1-q)\rho+q\text{tr}_{k}(\rho)\otimes|\bot\rangle\langle\bot|\,$, (3)
where with probability $q$ it is lost. Here, $\text{tr}_{k}(\rho)$ traces out the $k$th qubit. We assume loss occurs on all qubits equally with the total channel given by
@@
For a fully entangled state, loss on any qubit collapses the state into a mixture, and it cannot be used for further processing. As a simple example, let us consider the $N$-qubit GHZ state
@@
Loss of any qubit collapses the state into an (N1)(N-1) qubit mixed state
@@
the channel only traces out the $k$th qubit state $|\phi_{k}\rangle$, while the other N1N-1 qubits remain in a pure state. Thus, product states are more robust, which we will showcase in the following sections.
@@ III.1 Unencoded transfer
Now, let us consider a naive protocol to transport states across the qubit loss channel, which we will see has a low success probability. We perform a direct unencoded transfer of the $R$ copies. Here, we apply channel $\mathcal{E}(|\psi\rangle)$ on the pure $N$-qubit entangled state $|\psi\rangle$ and with a chance of $(1-q)^{N}$ no loss occurs on any qubit. The chance of successfully transferring the state has an exponentially small probability. Now we repeat the protocol with $R$ copies of $|\psi\rangle$. We define the failure probability $Q_{\text{e}}$ of the unencoded protocol as the probability that none of the $R$ transmitted copies of the entangled state $|\psi\rangle$ are received intact, i.e., every copy suffers a loss on at least one qubit. Equivalently, the protocol succeeds only if at least one of the $R$ copies survives with no qubit lost.
Hence, the failure probability of the unencoded protocol is
@@
which implies that we require exponentially large $R$ with $N$ for fixed failure probability $Q_{\text{e}}$.
Input: $R$ copies of entangled $n$-qubit state $|\psi\rangle$, disentangler $U$, qubit-loss channel $\mathcal{E}$
Output: Recovered state $|\tilde{\psi}\rangle$ or protocol failure
1
2for $i=1$ to $R$ do
3 Prepare new copy: $|\psi\rangle$;
4 Disentangle: $|\phi\rangle\leftarrow U|\psi\rangle$;
5 Transfer qubits: $\tilde{\rho}\leftarrow\mathcal{E}(|\phi\rangle\langle\phi|)$;
6   Apply leakage detection to each qubit in $\tilde{\rho}$ and keep accepted qubits (Fig. 3);
7
8 end for
9
10if any qubit index $j\in\{1,\dots,N\}$ has no surviving copy across $R$ trials then
11 Abort protocol;
12
13 else
14   Select one surviving copy for each qubit index $j$ to build product state $\bigotimes_{j=1}^{N}|\phi_{j}\rangle$;
15   Reconstruct: $|\tilde{\psi}\rangle\leftarrow U^{\dagger}\bigotimes_{j=1}^{N}|\phi_{j}\rangle$;
16
17 end if
Algorithm 1 DQAE protocol for qubit-loss channels
@@ III.2 DQAE
Now, we show that the DQAE can transport through qubit-loss channels with exponentially fewer copies than the unencoded scenario. Our DQAE algorithm is summarized in Alg. 1 as well as Fig. 2. We start by compressing the incoming $N$-qubit state $|\psi\rangle$ into a single-qubit product state using the DQAE $U$. Here, we assume that we already possess a pre-trained disentangler $U$. We are now left with an $N$ qubit product state of the form $U|\psi\rangle=\otimes_{j=1}^{N}|\phi_{j}\rangle$, where $|\phi_{j}\rangle\in\mathcal{H}^{2}$ is a single qubit state. Product states are robust against loss, as in the event of loss, the remaining qubits remain pure. Our protocol for product states is as follows: We have $R$ copies of $U|\psi\rangle$. After applying $\mathcal{E}(U|\psi\rangle)$, some qubits are lost; however, we assume that of the single qubit states $|\phi_{j}\rangle$, at least one of each has survived. One can check whether a qubit has survived with the leakage detection protocol [37] as shown in Fig.3, which also has been realized in experiment [38].
@@
Figure 3: Leakage detection protocol [37]. If data qubit $|\psi\rangle$ is present, measurement of the ancilla yields outcome 11 and leaves $|\psi\rangle$ unchanged. If the data qubit was lost (i.e. became an inaccessible orthogonal state $|\bot\rangle$), measurement of the ancilla yields outcome 0.
We now compute the failure probability of the product protocol. In this scenario, we define the failure probability $Q_{\text{p}}$ of the product protocol as the probability that for at least one qubit index $j\in\{1,\dots,N\}$, all $R$ copies of $|\phi_{j}\rangle$ are lost. In other words, the protocol fails unless each qubit has at least one surviving copy. For $N=1$, the probability that at least one qubit survives the channel is $1-q^{R}$. For $N$ qubits, the product protocol has a failure probability
@@ III.3 Fidelity
We now show how to lower bound the fidelity of the protocol of Fig. 2. We have an entangled state $|\psi\rangle$ and then run our protocol of Fig. 2. Let us assume that during the protocol, we send $N$ copies of disentangled $U|\psi\rangle$ through the loss channel. Further, we assume that for the $k$th copy, we lose all qubits except the $k$th qubit, such that the remaining state is given by $\rho_{k}=\text{tr}_{\bar{k}}(|\psi\rangle)$. Here, we define $\text{tr}_{\bar{k}}(.)$ as the partial trace over all qubits except the $k$th qubit. We can detect which qubits survived using the leakage detection of Fig. 3, and use the reconstruction step to assemble the $N$ remaining qubits $\rho_{k}$ into a single state $\rho=\bigotimes_{k}\rho_{k}$. If the disentangler is optimal, we can perfectly reconstruct the original state via $|\psi\rangle=U^{\dagger}\rho U$, as each $\rho_{k}$ is pure. However, what if the disentangler is not working perfectly? One can quantify the performance of the disentangler using the purity of the reduced density matrix of the $k$th qubit, which is given by
$E_{k}(|\psi\rangle=\text{tr}(\rho_{k}^{2})=\text{tr}(\text{tr}_{\bar{k}}(|\psi\rangle\langle\psi|)^{2})\,\,.$ (12)
When purity $E_{k}=1$, then the $k$th qubit is an unentangled single-qubit state. Thus, an optimal disentangler must have $E_{k}=1$ for all $k$.
If the disentangler does not optimally disentangle states, it creates an entangled state with $U|\psi\rangle$ with $E_{k}<1$ for at least one $k$. In this case, for the reconstruction step of Fig. 2, tracing out for each individual qubit $k$ results in a mixed state $\rho_{k}=\text{tr}_{\bar{k}}(U|\psi\rangle)$. The reconstructed state is then given by $\rho_{\text{f}}=U^{\dagger}\bigotimes_{k=1}^{N}\text{tr}_{\bar{k}}(U|\psi\rangle)U$. We can compute the fidelity between the original state $|\psi\rangle$ and the reconstructed state as
@@
where $F=1$ only when the DQAE encoded state $U|\psi\rangle$ is a product state, and is bounded by $2^{-N}\leq F\leq 1$.